OpenStudy (anonymous):

How do i solve this quadratic equation? (x^2-1)(x+2)=0

OpenStudy (anonymous):

HINT: x²-1 = (x+1)(x-1)

OpenStudy (anonymous):

do you have to expand it?

OpenStudy (anonymous):

No. You don't have to (thank god). Here, you have (x+1)(x-1)(x+2) = 0 Since this entire thing is equal to zero, we know that one of the factors must be zero. The only way to multiply and get zero is to multiply by zero. So x+1 = 0 or x - 1 = 0 or x+2 = 0 Apparently there is three possible answers for this problem. Does this make sense?

OpenStudy (anonymous):

oh i see so if theres three of them instead of the usual two then there's three possible answers?

OpenStudy (harsimran_hs4):

MIND IT : it`s not a quadratic equation. it`s cubic equation

OpenStudy (anonymous):

Most likely, unless you have two same factors.

OpenStudy (harsimran_hs4):

so it will have 3 answer

OpenStudy (anonymous):

and you know there is no same factors so three answers.

OpenStudy (anonymous):

Does this make sense? Yes it is cubic equation :D

OpenStudy (harsimran_hs4):

@RD⁴² so what if there are 2 same roots still there are 3 answer to cubic equation

OpenStudy (anonymous):

Let me show you a example. (x+1)(x+1)(x-3)=0 x = -1, 3 since it has two same factors, there is two answers for this equation. If all three factors are same, then one answer.

OpenStudy (anonymous):